# MMRA: A Benchmark for Multi-granularity Multi-image Relational Association

### v1

@@ 4 Result Analysis
Idefics2 58.11 32.10 65.71 74.16 28.30 48.89 55.56 90.00 38.89 72.22 86.00 49.21
Qwen-VL-Chat 42.11 35.80 59.05 60.67 33.96 44.44 31.75 55.56 27.78 57.78 49.00 14.29
Phi3v 51.84 49.38 66.67 62.92 48.11 58.89 46.03 85.56 34.44 56.67 51.00 20.64
IQ Mantis-Idefics2 57.90 35.80 62.86 70.79 37.74 51.11 42.86 81.11 26.67 74.44 91.00 56.35
@@
Table 2: The main results of current LVMLs and LLMs on our MMRM benchmark. The IQ and QO represent the Image+Question input and Question Only input, respectivelu.
@@ 4.1 Overall Analysis
As shown in Table 2, when inputting all question and image pairs (Image+Question setting), the close-source model(i.e., GPT-4v, GPT-4o, Gemini-Pro, and Gemini-Flash) achieve the best performance on our MMRA benchmark, with overall accuracy surpassing 60%. In contrast, the overall performance of other open-source multi-image LVLMs ranges from 50% to 60%, with the exception of Qwen-VL-Chat whose scores only 42.11%.
@@
It is worth mentioning that subtasks related to spatial perception (i.e., RelativePosition, NearSimilarEntity, Layout, and ObservationAngle) remain challenging for LVLMs, as most models achieve accuracy below 50% for these subtasks.
For the LLMs, both the question and the options are incorporated as inputs to evaluate them on benchmark (i.e., Question Only setting). The performance of LLMs on the ’UsageSimilarity’ task consistently exceeds 60%, which is comparable to the performance of multi-image LVLMs. This suggests that a significant portion of the reasoning required for the ’UsageSimilarity’ subtasks relies on commonsense knowledge inherent in the language model component of LVLMs.
Additionally, at the question only (QO) setting, we observe that all models, regardless of series or parameters size, demonstrate very similar performance across each task. This consistency indicates that our benchmark effectively mitigates textual information leakage and the accurate answer relies on visual information for accurately answering.
@@ 4.2 Impact of Image Input
As shown in Table 2, when provided with both image pairs and questions (i.e., Image+Question setting), the multi-image LVLMs demonstrate significantly better performance compared to LLMs. To better demonstrate the performance of LVLMs on various tasks, we calculate the average performance of all LLMs on each task as a basic performance of LLMs. By comparing this basic performance with the performance of LVLMs, we can quantify the actual improvement brought by incorporating images.
@@
As shown in the Fig 4, compared to the entity level, the relative improvement at the image level is better, which also indirectly confirms that the entity-level multi-image relation association task requires the model to be able to perceive more image details (the relative improvement at the entity level is around 0.1, while that of the image level is around 0.3). At the entity level, while the overall performance on the MentalitySimilarity subtask is comparable to other subtask, the improvement attributed to the inclusion of images is the most significant.
This suggests that the current LVLMs have developed a robust capacity to perceive mental states during pre-training. As a result, multi-image LVLMs can effectively harness the information in images to analyze relation between multiple images in the context of individuals’ mental states. This observation offers valuable insight into enhancing current LVLMs. The challenge related with fine-grained multi-image relation associating tasks stems from the fact that current multi-image LVLMs do not optimally utilize the detailed information within images. Instead, they tend to rely more on the language model component to process the textual information in the questions.
4.3 Impact of Image’s Description
@@
Table 3: The results of DQ and IDQ setting on our MMRM benchmark.
@@
We use the LLaVA-NeXT-100B to obtain the image caption and input it as the extra information. The results of DQ and IDQ settings are presented in the Tab 3. Under the DQ setting, with the combination of descriptions of image pair, all the LLMs’ performance are all highly improved, and the overall result of Qwen2-72B-Chat surpass the Gemini-Flash and is second only to GPT-4v, GPT-4o and Gemini-Pro. This also demonstrates that the multi-image understanding capability of LVMLs mainly stems from the content that they precept from image. As for the DQ setting, after including image descriptions, the performance of LVMLs don’t change significantly. This suggests that the image descriptions obtained by LLaVA-NeXT-100B overlap with the content perceived by LVMLs themselves. Furthermore, it also indicates that the bottleneck in enhancing the multi-image perception capabilities of LVLMs is the reasoning ability of the language model.
@@
As for the image level task, the LVLMs’ performance is not obviously changed at IDQ setting, while the LLMs’ results are close to that of VLMs with the input of images’ descriptions. It demonstrates that the multi-image perception perception at image level relies on the visual module of LVMLs. With regarding to the tasks at entity level, in the IDQ setting, the performance of LVMLs varied the most on the MentalitySimilarity (MS) task, even surpassing GPT-4v and GPT-4o. This indicates that associating the relations of mental states among creatures in multiple images requires highly advanced capabilities from both the language model and the image model
@@ 4.4 Error Analysis
To better analyze the shortcomings of LVLMs, we examined instances where GPT-4o made errors on relatively challenging subtasks such as RelativePosition, MadeOf, Nearby, and Layout.
As presented in Fig 5, LVLMs often select entities that do not appear in the image when answering fine-grained questions. For example, for subtasks like ’RelativePosition’ and ’Nearby’, LVLMs sometimes choose options featuring entities that are not present in the image (e.g., beer and tray).
@@ 5 Conclusion
6 Related Work
Large Visual Language Models.
With the emergence of Large Language Models(LLMs), researchers have applied it to the multimodal perception field. In recent years, more and more Large Visual Language Models (LVLMs) have achieved excellent success in combining vision knowledge with language on single-image task. BLIP2 [14] uses a Q-Former to align the visual knowledge with the textual information during the pre-training phase. LLaVA [16] is pre-trained on the GPT-4 generated multimodal language-image instruction-following data and uses a projection to fuse the vision module and the LM. LLaVA-Next [15] improves the single-image performance with a cost of an increasing number of image tokens. Based on the framework LLaVA mode, QwenVL [3], CogVLM[27], and Yi-VL[1] all achieve significant success through huge number of pre-training data. Those LVLMs all demonstrate exceptional reasoning ability on single image tasks, such as TextVQA [23], VQAV2 [8], MMBench[28], GQA[9], etc. Although Fuyu-8B [^1] https://www.adept.ai/blog/fuyu-8b, Kosmos2 [17], and Flamingo [2] support the interleaved image-text input, they don’t optimize in the multi-image reasoning.
Multi-Image Perception Model and Task.
Currently, some researchers have realized the importance of the multi-image ability of LVLMs. Excepting Kosmos2, Fuyu and Flamingo, There are some models which support input multi images (e.g. Mantis, Idefic2, Phi3v and Mantis-Idefic2 [25, 13, 18, 11]). Besides, the Emu2[25] is the generative multimodal model that supports the interleaved text-image inputs. And the video understanding (e.g., Video-LLaMA [29]) also could input the multi-image inputs, but it is relatively worse than the VLMs. Meanwhile, there is also a lack of comprehensive and systematic evaluation of multi-image LVLMs. The difference description is the first task and researchers have developed many datasets, such as Spot-the-Diff and Birds-to-Words [10], etc. However, they are all generative tasks. Recently, the MuirBench [26] and the multi-image under standing benchmark [30] focus on evaluating the LVMLs’ ability, but they don’t systematically define the relation among images in real-life scenario.
Commonsense Reasoning.
@@ Appendix A Result Exact Matching Rule
For Multiple-Choice questions: First, we use regular expressions to attempt to directly extract the matching content within parentheses, i.e., extracting Answer: ”A” from ”(A)”. If this is unsuccessful, we then attempt to match option labels (A-D) from the entire response content and return the option with the highest match count. If the response does not contain any option label information, we try to match the option content directly within the response and return the corresponding option label.
For True/False questions: We use regular expressions to match ”yes” or ”no” within the response content. If there are multiple matches, we return the result that appears the most frequently.
@@ Appendix B Designed Template
Table 4: The designed prompt template for the task in our MMRM benchmark.
@@ Appendix C Sampled examples from MMRA benchmark
In order to comprehensively show our benchmark, we select a sample for each task and present then in the Tab 6. We design two kind of task (i.e., Choice Question and T/F Question). For each example, we show the image pair, question and options.

### v2

@@ 4 Result Analysis
Idefics2 56.93 37.04 65.66 69.66 28.30 44.44 53.97 87.78 36.67 72.22 88.00 45.24
Mantis-Idefics2 57.59 35.80 62.63 68.54 41.51 52.22 41.27 82.22 20.00 74.44 91.00 56.35
Phi3v 51.75 48.15 64.65 62.92 47.17 61.11 46.03 86.67 34.44 56.67 51.00 20.63
IQ Qwen-VL-Chat 47.45 37.04 58.59 68.54 34.91 48.89 41.27 73.33 33.33 61.11 50.00 23.02
@@
Table 2: The main results of current LVMLs and LLMs on our MMRA benchmark. The IQ and QO represent the Image+Question input and Question Only input, respectively.
@@ 4.1 Overall Analysis
As shown in Table 2, when inputting all question and image pairs (Image+Question setting), the close-source model(i.e., GPT-4v, GPT-4o, Gemini-Pro, and Gemini-Flash) achieves the best performance on our MMRA benchmark, with overall accuracy surpassing 60%. In contrast, the overall performance of other open-source multi-image LVLMs ranges from 50% to 60%, with the exception of Qwen-VL-Chat whose score is only 47.45%.
@@
It is worth mentioning that subtasks related to spatial perception (i.e., RelativePosition, NearSameEntity, Layout, and ObservationAngle) remain challenging for LVLMs, as most models achieve accuracy below 50% for these subtasks.
For the LLMs, both the question and the options are incorporated as inputs to evaluate them on the benchmark (i.e., Question Only setting). The performance of LLMs on the ’UsageSimilarity’ task consistently exceeds 60%, which is comparable to the performance of multi-image LVLMs. This suggests that a significant portion of the reasoning required for the ’UsageSimilarity’ subtasks relies on commonsense knowledge inherent in the language model component of LVLMs.
Additionally, at the question-only (QO) setting, we observe that all models, regardless of series or parameter size, demonstrate very similar performance across each task. This consistency indicates that our benchmark effectively mitigates textual information leakage and those subtasks rely on visual information for accurately answering.
@@ 4.2 Impact of Image Input
As shown in Table 2, when provided with both image pairs and questions (i.e., the Image+Question setting), multi-image LVLMs demonstrate significantly better performance compared to LLMs under the QO setting (i.e., Question Only). To highlight the performance improvement of LVLMs due to image input across various tasks, we calculate the average performance of all LLMs on each task as a baseline. By comparing this baseline with the performance of LVLMs, we can quantify the actual enhancement brought about by incorporating images.
@@
As shown in the Fig 5, compared to the entity level, the relative improvement at the image level is better, which also indirectly confirms that the entity-level multi-image relation association task requires the model to be able to perceive more image details (the relative improvement at the entity level is around 0.1, while that of the image level is around 0.3). At the entity level, while the overall performance on the MentalitySimilarity subtask is comparable to other subtasks, the improvement attributed to the inclusion of images is the most significant. This suggests that the current LVLMs have developed a robust capacity to perceive mental states during pre-training. As a result, multi-image LVLMs can effectively harness the information in images to analyze the relation between multiple images in the context of individuals’ mental states. This observation offers valuable insight into enhancing current LVLMs. The challenge related to fine-grained multi-image relation association tasks stems from the fact that current multi-image LVLMs do not optimally utilize the detailed information within images. Instead, they tend to rely more on the language model component to process the textual information in the questions.
4.3 Impact of Image Descriptions
@@
Table 3: The results of DQ and IDQ setting on our MMRA benchmark.
@@
We use the LLaVA-NeXT-100B to obtain the image caption and input it as the extra information. The results of DQ and IDQ settings are presented in the Tab 3. Under the DQ setting, with the combination of descriptions of image pair, all the LLMs’ performance is highly improved, and the overall result of Qwen2-72B-Chat surpasses the Gemini-Flash and is second only to GPT-4v, GPT-4o, and Gemini-Pro. This also demonstrates that the multi-image understanding capability of LVLMs mainly stems from the content that they precept from images. As for the DQ setting, after including image descriptions, the performance of LVLMs doesn’t change significantly. This suggests that the image descriptions obtained by LLaVA-NeXT-100B overlap with the content perceived by LVLMs themselves. Furthermore, it also indicates that the bottleneck in enhancing the multi-image perception capabilities of LVLMs is the reasoning ability of the language model.
@@
As for the image level task, the LVLMs’ performance is not obviously changed at IDQ setting, while the LLMs’ results are close to that of VLMs with the input of images’ descriptions. It demonstrates that the multi-image perception at the image level relies on the visual module of LVLMs. With regard to the tasks at the entity level, in the IDQ setting, the performance of LVLMs varied the most on the MentalitySimilarity (MS) task, even surpassing GPT-4v and GPT-4o. This indicates that associating the relations of mental states among creatures in multiple images requires highly advanced capabilities from both the language model to process the information obtained from the image.
4.4 Image Sequence Perception Ability
Understanding the sequential order of images is crucial for interpreting the relations between multiple images. The ability of a model to comprehend image sequences is essential for tackling complex multi-image tasks, such as sorting images. In certain subtasks of the MMRA Benchmark, the sequence of input images can influence the response to the associated questions. For instance, in the SimilarMaterial subtask, some options describe entities present in both image1 and image2. Altering the sequence of these input images could lead to a scenario where a correct answer is no longer available.
To examine the ability to perceive image sequences, we adjusted the input image sequence for four specific subtasks: Relative Position, Similar Material, Near Similar Entity, and Observation Angle. Each subtask has options that are directly related to the image sequence. Additionally, we introduced a new option, “All of the above options are incorrect” as the correct choice. We then evaluated the performance of LVLMs on these subtasks.
As illustrated in Table 4, we present the accuracy metrics of various LVLMs. Idefics2 demonstrates commendable image sequence perception, achieving an overall score close to 60%. In contrast, most current LVLMs exhibit inadequate image sequence perception abilities, with overall scores below 35%. This discrepancy suggests that current open-source LVLMs have not adequately addressed image sequence tasks during their pre-training processes.
Model Overall RP SM NSE OA
Idefics2 59.94 65.43 78.30 82.54 13.49
Mantis 0.00   0.00   0.00   0.00   0.00
Phi3v 33.20 41.98 47.17 31.75 11.90
Qwen-VL 0.63   0.00   0.94   1.59   0.00
Table 4: The results of the Sequence Perception task result.
@@ 4.5 Error Analysis
To better analyze the shortcomings of LVLMs, we examined instances where GPT-4o made errors on relatively challenging subtasks such as RelativePosition, MadeOf, NearSameEntity, and Layout.
As presented in Fig 6, LVLMs often select entities that do not appear in the image when answering fine-grained questions. For example, for subtasks like ’RelativePosition’ and ’NearSameEntity’, LVLMs sometimes choose options featuring entities that are not present in the image (e.g., beer and tray).
@@
5 Related Work
Large Visual Language Models.
With the emergence of Large Language Models(LLMs), researchers have applied it to the multimodal perception field. In recent years, more and more Large Visual Language Models (LVLMs) have achieved excellent success in combining vision knowledge with language on single-image tasks. BLIP2 [13] uses a Q-Former to align the visual knowledge with the textual information during the pre-training phase. LLaVA [16] is pre-trained on the GPT-4 generated multimodal language-image instruction-following data and uses a projection to fuse the vision module and the LM. LLaVA-Next [15] improves the single-image performance with a cost of an increasing number of image tokens. Based on the framework LLaVA mode, QwenVL [3], CogVLM[28], and Yi-VL[1] all achieve significant success through a huge number of pre-training data. Those LVLMs all demonstrate exceptional reasoning ability on single image tasks, such as TextVQA [24], VQAV2 [7], MMBench[30], GQA[8], etc. Although Fuyu-8B [^1] https://www.adept.ai/blog/fuyu-8b, Kosmos2 [17], and Flamingo [2] support the interleaved image-text input, they don’t optimize in the multi-image reasoning.
Multi-Image Perception Model and Task.
Currently, some researchers have realized the importance of the multi-image ability of LVLMs. Excepting Kosmos2, Fuyu and Flamingo, There are some models which support input multi images (e.g. Mantis, Idefic2, Phi3v and Mantis-Idefic2 [26, 12, 18, 10]). Besides, the Emu2[26] is the generative multimodal model that supports the interleaved text-image inputs. And the video understanding models [32, 20] also have the multi-image perception ability and the Video-LLaMA [32] also support the multi-image inputs, but it is relatively worse than the VLMs. Meanwhile, there is also a lack of comprehensive and systematic evaluation of multi-image LVLMs. The difference description is the first task and researchers have developed many datasets, such as Spot-the-Diff and Birds-to-Words [9], etc. However, they are all generative tasks. Recently, the MuirBench [27] and the multi-image under standing benchmark [33] focus on evaluating the LVLMs’ ability, but they don’t systematically define the relation among images in real-life scenario.
Commonsense Reasoning.
@@ Appendix A Result Exact Matching Rule
For Multiple-Choice questions: First, we use regular expressions to attempt to directly extract the matching content within parentheses, i.e., extracting Answer: “A” from “(A)”. If this is unsuccessful, we then attempt to match option labels (A-D) from the entire response content and return the option with the highest match count. If the response does not contain any option label information, we try to match the option content directly within the response and return the corresponding option label.
For True/False questions: We use regular expressions to match “yes” or “no” within the response content. If there are multiple matches, we return the result that appears the most frequently.
@@ Appendix B Designed Template
Table 5: The designed prompt template for the task in our MMRA benchmark.
@@ Appendix C Sampled examples from MMRA benchmark
In order to comprehensively show our benchmark, we select a sample for each task and present then in the Tab 7. We design two kinds of tasks (i.e., Choice Question and T/F Question). For each example, we show the image pair, question and options.